# Fast post-process Bayesian inference with Variational Sparse Bayesian Quadrature

### v2

@@ 3.3 Noise shaping
Figure 3: Illustration of noise shaping effect. Noise shaping improves the selection of inducing points and the approximation of a sparse GP in the high posterior density region. A better sparse GP consequently leads to an improved variational posterior. The red triangles are locations of inducing points selected via Eq. 24. The orange line represents the sparse GP prediction. The green dashed line represents the variational posterior. (a) True density function (black line) and the observations (black circles). (b) Sparse GP approximation to the log-density function without noise shaping. The black curve is the ground truth log-density function. (c) Variational posterior without noise shaping. (d) Noise shaping curve corresponding to Eq. 26. (e) Sparse GP approximation with noise shaping. (f) Variational posterior with noise shaping.
@@
$\sigma_{\text{shape}}(\Delta y)=\exp((1-\rho)\log\sigma_{\text{min}}+\rho\log\sigma_{\text{med}})+\mathbf{1}_{\Delta y\geq\theta_{\sigma}}{\lambda_{\sigma}}(\Delta y-\theta_{\sigma}),$ (26)
@@
$\log\tilde{p}(\mathbf{y}\mid\mathbf{f})=\sum_{n=1}^{N}\log\tilde{p}(y_{n}|f_{n})=C-\sum_{n=1}^{N}\frac{(y_{n}-f(\mathbf{x}_{n}))^{2}}{2\sigma^{2}_{\text{tot}}(\mathbf{x}_{n},y_{n})},$ (27)
@@
$\mathbb{E}_{q(\mathbf{f})}[\log\tilde{p}(\mathbf{y}\mid\mathbf{f})]=\text{const}+\sum_{n=1}^{N}w_{n}\mathbb{E}_{q(f_{n})}[\log p(y_{n}|f_{n})],\quad w_{n}\equiv\frac{\sigma^{2}_{\text{obs}}(\mathbf{x}_{n})}{\sigma^{2}_{\text{tot}}(\mathbf{x}_{n},y_{n})}\leq 1.$ (29)
That is, by assigning larger ‘shaping’ noise to low-density observations we are downweighing their role in the sparse GP representation, guiding the sparse GP to better represent higher-density regions. Noise shaping also helps during inducing point selection (Eq. 24), as shown in Figure 3.
@@ 3.4 Approximation error
$\|q_{\phi}-q_{\phi_{e}}\|_{\text{TV}}<\sqrt{\gamma_{1}/2}+\sqrt{\gamma_{2}/2}+(1-\mathrm{exp}(-K_{l}\sqrt{\gamma/2})).$ (31)
@@ 4.1 Baseline methods
As few methods afford post-process and black-box inference of the posterior, we compare vsbq again two baselines: neural network regression (nnr) and the popular Laplace approximation (laplace).
@@
For a direct comparison, we develop a post-process inference algorithm based on a deep neural network regression surrogate (nnr) instead of a sparse GP surrogate, otherwise leaving the post-process procedure (Figure 2) as much as possible the same. This is a competitive baseline since deep neural networks exhibit strong regression performance in the presence of a large number of data (Goodfellow et al., 2016). For the network architecture, we use a multilayer perceptron (MLP) with an input layer of dimension $D$, four hidden layers of $1024$ units, and an output layer for predicting the log-density value. The activation function is chosen to be the rectified linear units (ReLu; Goodfellow et al., 2016, Chapter 6). In addition, we add a negative quadratic mean function to the neural network output to ensure that it is a valid log-density surrogate function. [^6] An MLP with ReLu activations is a continuous piecewise affine function (Mukherjee et al., 2018), and therefore adding a trainable negative quadratic mean function ensures the integrability of the exponentiated surrogate. The negative quadratic mean function is the same as the one used for the (sparse) GP (see Eq. 10). In total, the surrogate function $g$ is:
@@
where $\sigma_{\text{tot}}^{2}(\mathbf{x}_{n},y_{n})=\sigma_{\text{obs}}^{2}(\mathbf{x}_{n})+\sigma_{\text{shape}}^{2}(\Delta y_{n})$, as per noise shaping. [^7] We found that noise shaping empirically helped stabilize the neural network training. We optimize the neural network parameters by minimizing the objective in Eq. 33 with AdamW optimizer (Loshchilov and Hutter, 2019). For regularization, we considered the ‘weight decay’ hyperparameter $\alpha\in\{0,0.01,0.1\}$, selecting the best neural network surrogate based on the loss on a split validation dataset. Finally, we use stochastic variational inference with the reparametrization trick (Kingma and Welling, 2013) to compute the approximate posterior. This part is the same as vsbq except that the expected log joint of the surrogate in Eq. 2 is approximated via Monte Carlo samples rather than calculated exactly via sparse Bayesian quadrature. For more details, see Supplementary Material C.3.
@@
The Laplace approximation method (laplace) computes a multivariate normal approximation of the posterior centered at the MAP location in the unbounded parameter space (see “Inference space” in Section 4.2 below), providing both a posterior approximation and an estimate of the marginal likelihood (MacKay, 2003). Despite its simplicity, the Laplace approximation is often used in practice for its efficiency and can yield reasonable posterior approximations (Piray et al., 2019; Daxberger et al., 2021). Note that laplace requires additional log-density evaluations for numerically estimating the Hessian and is not easily applicable to noisy evaluations, so this is not a fully post-process inference method, but we include it as a popular baseline and useful performance reference.
@@ 4.2 Procedure and metrics
vsbq, nnr, and laplace all operate in an unbounded parameter space also known as inference space. The unbounded inference space is necessary to define and manipulate the multivariate normals (and mixtures thereof) used by all our algorithms. Parameters that are originally subject to bound constraints are mapped to the inference space via a shifted and rescaled probit transform, with an appropriate Jacobian correction to the log-density values. A similar approach is common in probabilistic inference software (Carpenter et al., 2017; Acerbi, 2018). Solutions of vsbq, nnr, and laplace are transformed back to the original space via the corresponding inverse transform, for computing the metrics against the ground truth posterior.
@@
The optimization trace points and corresponding log-density values are collected as the training dataset for vsbq and nnr. For each problem, we repeated the entire optimization procedure – each with multiple MAP estimation runs, as explained above – ten times with different random seeds. This yielded ten different training sets per problem, used to assess the robustness and reliability of the methods. The number of inducing points for the sparse GP is set to $100D$. In the case of a noisy target, we further vary the observation noise level to study the noise sensitivity of vsbq and nnr. For the Laplace approximation, we first find the MAP point by transforming the parameter space to unbounded if needed, via a nonlinear mapping. [^8] For the Laplace approximation to be valid, it is necessary to find the MAP estimate in the unbounded space subsequently used to compute the multivariate normal approximation; this is particularly important if the mode is close to the bounds. We subsequently compute the Hessian matrix via adaptive numerical differentiation (Brodtkorb and D’Errico, 2022). Finally, we compute the performance metrics as detailed below.
@@
    $\text{GsKL}(p,q)=\frac{1}{2D}\left[D_{\text{KL}}\left(\mathcal{N}[p]||\mathcal{N}[q]\right)+D_{\text{KL}}(\mathcal{N}[q]||\mathcal{N}[p])\right],$ (35)
@@ 4.3 Synthetic problems
$p(\mathbf{x})$ $=-\frac{1}{2}\left(\frac{r-1/\sqrt{2}}{0.1}\right)^{2}+\log(\exp(\kappa x_{1}/r)/3+2\exp(-\kappa x_{1}/r)/3),$ (36)
@@
Figure 4: Two Moons. The black density contours and points denote ground truth samples. The orange density contours and points represent the posterior samples from (a) vsbq; (b) nnr; and (c) laplace. Both vsbq and nnr perfectly recover the bimodal density.
As shown in Figure 4 and Table 1, both vsbq and nnr reconstruct the bimodal target almost perfectly. By contrast, the Laplace approximation can only cope with unimodal posteriors and thus unsurprisingly fails in this case, despite its otherwise relative simplicity.
@@
$p(\mathbf{x})\propto e^{\mathcal{R}(x_{1},x_{2})}e^{\mathcal{R}(x_{3},x_{4})}\mathcal{N}(x_{5},x_{6};\mathbf{0},\mathbb{I})\mathcal{N}(\mathbf{x};\mathbf{0},3^{2}\mathbb{I}),$ (37)
@@
As shown in Table 2, both vsbq and nnr approximate this complex posterior well, with vsbq performing slightly better than nnr in terms of metrics. laplace does not give a satisfactory approximation either, due to the heavily non-Gaussian nature of the underlying posterior. A visualization of the approximate posteriors and the ground-truth posterior is provided in Supplementary Material C.5.
@@ 4.4 Real-world models
We consider a popular Bayesian observer model of time perception from cognitive neuroscience (Jazayeri and Shadlen, 2010; Acerbi et al., 2012; Acerbi, 2020). The key premise of Bayesian observer modeling in perception is that the experimental subject – which is the system being modeled – is herself performing Bayesian inference over the sensory stimuli, and employs Bayesian decision theory to report their perception (Pouget et al., 2013; Ma et al., 2023).
@@
In this specific sensorimotor timing experiment, in each trial human subjects had to reproduce the time interval $\tau$ between a mouse click and a screen flash, with $\tau\sim\text{Uniform}$[0.6, 0.975] s (Acerbi et al., 2012). We assume subjects had only access to a noisy sensory measurement $t_{\text{s}}\sim\mathcal{N}\left(\tau,w_{\text{s}}^{2}\tau^{2}\right)$, and their reproduced time $t_{\text{m}}$ was affected by motor noise, $t_{\text{m}}\sim\mathcal{N}\left(\tau_{\star},w_{m}^{2}\tau_{\star}^{2}\right)$, where $w_{\text{s}}$ and $w_{\text{m}}$ are Weber’s fractions, a psychophysical measure of perceptual and motor variability. We assume subjects estimated $\tau_{\star}$ by combining their sensory likelihood with an approximate Gaussian prior over time intervals, $\mathcal{N}\left({\tau};{\mu_{\text{p}}},{\sigma^{2}_{\text{p}}}\right)$, and took the mean of the resulting Bayesian posterior. For each trial we also consider a probability $\lambda$ of a ‘lapse’ (e.g., a misclick) producing a response $t_{\text{m}}\sim\text{Uniform}$[0, 2] s. Model parameters are $\bm{\theta}=(w_{\text{s}},w_{\text{m}},\mu_{\text{p}},\sigma_{\text{p}},\lambda)$, so $D=5$. We infer the posterior of a representative subject using published data from Acerbi et al. (2012).
@@
As shown in Figure 5, vsbq consistently gives a good posterior approximation across different levels of log-likelihood evaluation noise. In this case, even with a large noise $\sigma_{\text{obs}}=7$, vsbq surpasses the performance of laplace. Note that laplace is computed based on noiseless target evaluations since it does not support noisy evaluations. In this problem, nnr is considerably less robust, with several failed runs, and performs generally worse compared to vsbq. Plots of approximate posteriors and the ground-truth posteriors are visualized in Supplementary Material C.5.
@@
Figure 5: Bayesian timing model. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and black dashed horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). vsbq performs well across all noise levels and can outperform (noiseless) laplace even under high log-likelihood noise, whereas nnr demonstrates less robustness with several failed runs.
@@
Perceptual causal inference – inferring whether two sensory cues have the same common source – comprises a variety of models and tasks of major interest in computational and cognitive neuroscience (Körding et al., 2007; Cao et al., 2019; Shams and Beierholm, 2022). Here we consider a visuo-vestibular causal inference experiment representative of this class of models (Acerbi et al., 2018; Acerbi, 2020). In this experiment, participants were seated in a moving chair and asked to determine whether the direction of their movement ($s_{\text{vest}}$) corresponded to the direction of a looming visual field ($s_{\text{vis}}$) on a trial by trial basis. It is assumed that the subjects can only access noisy sensory measurements, denoted as $z_{\text{vest}}\sim\mathcal{N}\left(s_{\text{vest}},\sigma^{2}_{\text{vest}}\right)$ for vestibular information and $z_{\text{vis}}\sim\mathcal{N}\left(s_{\text{vis}},\sigma^{2}_{\text{vis}}(c)\right)$ for visual information. Here, $\sigma_{\text{vest}}$ represents the vestibular noise, while $\sigma_{\text{vis}}(c)$ represents the visual noise, with $c$ being one of three distinct levels of visual coherence ($c_{\text{low}},c_{\text{med}},c_{\text{high}}$) used in the experiment. To model the subjects’ responses, we use a heuristic ‘Fixed’ rule, which determines the source to be the same if the absolute difference between the visual and vestibular measurements is less than a threshold $\kappa$, i.e., $|z_{\text{vis}}-z_{\text{vest}}|<\kappa$. Additionally, the model incorporates a probability $\lambda$ of the subject providing a random response (Acerbi et al., 2018). The model parameters are $\theta=(\sigma_{\text{vis}}(c_{\text{low}}),\sigma_{\text{vis}}(c_{\text{med}}),\sigma_{\text{vis}}(c_{\text{high}}),\sigma_{\text{vest}},\lambda,\kappa)$, with a total of $D=6$ parameters. Here we fit data from subject S0 of Acerbi et al. (2018).
@@
Figure 6: Multisensory causal inference model. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and black dashed horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). vsbq performs well across low to moderate noise levels and outperforms (noiseless) laplace even at high noise levels. nnr also demonstrates strong performance across all noise levels.
@@
The performance metrics of all tested methods are plotted in Figure 6, as a function of log-likelihood observation noise. vsbq consistently outputs a good posterior approximation and only exceeds the metrics thresholds at large observation noise. laplace works reasonably for the noiseless target, still worse than vsbq, and slightly above the thresholds for MMTV and GsKL. In this problem, nnr performs well, slightly outperforming vsbq in some metrics in a higher observation noise regime. We visualize the approximate and ground-truth posteriors in Supplementary Material C.5.

### v3

@@ 3.2 Inducing points and hyperparameter selection
[…]
where $\mathbf{Q}_{\mathbf{X},\mathbf{X}}\equiv\mathbf{K}_{\mathbf{X},\mathbf{Z}}\mathbf{K}_{\mathbf{Z},\mathbf{Z}}^{-1}\mathbf{K}_{\mathbf{Z},\mathbf{X}}$. Note that the GP-elbo depends on the location of inducing points $\mathbf{Z}$ and GP hyperparameters $\bm{\xi}$, and consists of the log probability density function of a multivariate normal term, minus the trace of an error term.
@@ 3.3 Noise shaping
Figure 3: Illustration of noise shaping effect. Noise shaping improves the selection of inducing points and the approximation of a sparse GP in the high posterior density region. A better sparse GP consequently leads to an improved variational posterior. The red triangles are locations of inducing points selected via Eq. 25. The orange line represents the sparse GP prediction. The green dashed line represents the variational posterior. (a) True density function (black line) and the observations (black circles). (b) Sparse GP approximation to the log-density function without noise shaping. The black curve is the ground truth log-density function. (c) Variational posterior without noise shaping. (d) Noise shaping curve corresponding to Eq. 3.3. (e) Sparse GP approximation with noise shaping. (f) Variational posterior with noise shaping.
@@
$\sigma_{\text{shape}}(\Delta y)=$ $\exp((1-\rho)\log\sigma_{\text{min}}+\rho\log\sigma_{\text{med}})$
$+\mathbf{1}_{\Delta y\geq\theta_{\sigma}}{\lambda_{\sigma}}(\Delta y-\theta_{\sigma}),$ (27)
@@
$\begin{split}\log\tilde{p}(\mathbf{y}\mid\mathbf{f})&=\sum_{n=1}^{N}\log\tilde{p}(y_{n}|f_{n})\\ &=C-\sum_{n=1}^{N}\frac{(y_{n}-f(\mathbf{x}_{n}))^{2}}{2\sigma^{2}_{\text{tot}}(\mathbf{x}_{n},y_{n})},\end{split}$ (28)
@@
$\begin{split}&\mathbb{E}_{q(\mathbf{f})}[\log\tilde{p}(\mathbf{y}\mid\mathbf{f})]\\ &=\text{const}+\sum_{n=1}^{N}w_{n}\mathbb{E}_{q(f_{n})}[\log p(y_{n}|f_{n})],\end{split}$ (30)
where $w_{n}\equiv\frac{\sigma^{2}_{\text{obs}}(\mathbf{x}_{n})}{\sigma^{2}_{\text{tot}}(\mathbf{x}_{n},y_{n})}\leq 1$. That is, by assigning larger ‘shaping’ noise to low-density observations we are downweighing their role in the sparse GP representation, guiding the sparse GP to better represent higher-density regions. Noise shaping also helps during inducing point selection (Eq. 25), as shown in Figure 3.
@@ 3.4 Approximation error
$\begin{split}\|q_{\phi}-q_{\phi_{e}}\|_{\text{TV}}<&\sqrt{\gamma_{1}/2}+\sqrt{\gamma_{2}/2}\\ &+(1-\mathrm{exp}(-K_{l}\sqrt{\gamma/2})).\end{split}$ (32)
@@ 4.1 Baseline methods
We recall that with black-box inference we mean that the target lacks gradients and may be expensive to evaluate and possibly noisy. As few other methods afford post-process and black-box inference of the posterior, we compare vsbq again three baselines: black-box variational inference (bbvi), neural network regression (nnr) and the popular Laplace approximation (laplace). Of these, only nnr is also a post-process method. Further implementation details of vsbq are provided in Supplementary Material C.1.
Black-box variational inference.
When the target posterior density is a black-box, the reparametrization trick (Kingma and Welling, 2013) cannot be applied to estimate the elbo gradient for stochastic variational inference. Instead, one needs to resort to other techniques for computing the gradient of the elbo, such as the score function estimator, also known as the REINFORCE estimator (Ranganath et al., 2014), which often has higher variance compared to the reparametrization trick (Gal, 2016; Xu et al., 2019). Specifically, by differentiating Eq. 2, the elbo gradient $\nabla_{\bm{\phi}}\textsc{elbo}\left[q_{\bm{\phi}}\right]$ can be written as,
$\begin{split}&\nabla_{\bm{\phi}}\mathbb{E}_{\bm{\phi}}\left[\log\frac{p(\mathcal{D}|\mathbf{x})p(\mathbf{x})}{q_{\bm{\phi}}(\mathbf{x})}\right]\\ &=\mathbb{E}_{\bm{\phi}}\left[\nabla_{\bm{\phi}}\log q_{\bm{\phi}}(\mathbf{x})\left(\log p(\mathcal{D}|\mathbf{x})p(\mathbf{x})-\log q_{\bm{\phi}}(\mathbf{x})\right)\right].\\ \end{split}$ (33)
In addition, we leverage the control variates technique to reduce the gradient variance (see the supplement for details). We experiment with the following variational distributions for bbvi: a Gaussian with diagonal covariance matrix, a Gaussian with full-rank covariance matrix, and a mixture of Gaussians with $K=5$ and $K=50$ components, respectively, where the mixture of Gaussians (MoG) admits the same form as the variational posterior in vsbq. The Adam optimizer (Kingma and Ba, 2014) is used for optimizing the elbo with stochastic gradients. For each variational distribution choice, we applied grid search on the learning rate hyperparameter in $\{0.01,0.001\}$ and the number of Monte Carlo samples for gradient estimation in $\{1,10,100\}$ and reported the best result, according to the estimated elbo value. Since bbvi typically requires a large number of target density evaluations for convergence, in all the experiments we assign ten times more evaluation budget to bbvi than vsbq to make it a stronger baseline for reference. Moreover, note that this is not a post-process method, but we include it as a reasonable performance reference. For more details, see Supplementary Material C.3.
@@
For a direct comparison with our method, we develop a post-process inference algorithm based on a deep neural network regression surrogate (nnr) instead of a sparse GP surrogate, otherwise leaving the post-process procedure (Figure 2) as much as possible the same. This is a competitive baseline since deep neural networks exhibit strong regression performance in the presence of a large number of data points (Goodfellow et al., 2016). For the network architecture, we use a multilayer perceptron (MLP) with an input layer of dimension $D$, four hidden layers of $1024$ units, and an output layer for predicting the log-density value. The activation function is chosen to be the rectified linear units (ReLu; Goodfellow et al., 2016, Chapter 6). In addition, we add a negative quadratic mean function to the neural network output to ensure that it represents a valid log-density surrogate function. [^7] An MLP with ReLu activations is a continuous piecewise affine function (Mukherjee et al., 2018), and therefore adding a trainable negative quadratic mean function ensures the integrability of the exponentiated surrogate. The negative quadratic mean function is the same as the one used for the (sparse) GP (see Eq. 17). In total, the surrogate function $g$ is:
@@
where $\sigma_{\text{tot}}^{2}(\mathbf{x}_{n},y_{n})=\sigma_{\text{obs}}^{2}(\mathbf{x}_{n})+\sigma_{\text{shape}}^{2}(\Delta y_{n})$, as per noise shaping. [^8] We found that noise shaping empirically helped stabilize the neural network training. We optimize the neural network parameters by minimizing the objective in Eq. 35 with the AdamW optimizer (Loshchilov and Hutter, 2019). For regularization, we considered the ‘weight decay’ hyperparameter $\alpha\in\{0,0.01,0.1\}$, selecting the best neural network surrogate based on the loss on a split validation dataset. Finally, we use stochastic variational inference with the reparametrization trick (Kingma and Welling, 2013) to compute the approximate posterior. This part is the same as vsbq except that the expected log joint of the surrogate in Eq. 2 is approximated via Monte Carlo samples rather than calculated exactly via sparse Bayesian quadrature. For more details, see Supplementary Material C.4.
@@
The Laplace approximation method (laplace) computes a multivariate normal approximation of the posterior centered at the MAP location in the unbounded parameter space (see “Inference space” in Section 4.2 below), providing both a posterior approximation and an estimate of the marginal likelihood (MacKay, 2003). Despite its simplicity, the Laplace approximation is often used in practice for its efficiency and can yield reasonable posterior approximations (Piray et al., 2019; Daxberger et al., 2021). Note that laplace requires additional log-density evaluations for numerically estimating the Hessian and is not easily applicable to noisy evaluations, so this is not a fully post-process inference method, but we also include it as a popular baseline and performance reference.
@@ 4.2 Procedure and metrics
vsbq, nnr, laplace and bbvi all operate in an unbounded parameter space also known as inference space. The unbounded inference space is necessary to define and manipulate the multivariate normals (and mixtures thereof) used by all our algorithms. Parameters that are originally subject to bound constraints are mapped to the inference space via a shifted and rescaled probit transform, with an appropriate Jacobian correction to the log-density values. A similar approach is common in probabilistic inference software (Carpenter et al., 2017; Acerbi, 2018). The approximate posteriors are transformed back to the original space via the corresponding inverse transform, for computing the metrics against the ground truth posterior.
@@
The optimization trace points and corresponding log-density values are collected as the training dataset for vsbq and nnr. For each problem, we repeated the entire optimization procedure – each with multiple MAP estimation runs, as explained above – ten times with different random seeds. This yielded ten different training sets per problem, used to assess the robustness and reliability of the methods. The number of inducing points for the sparse GP is set to $100D$. For bbvi, as mentioned in Section 4.1, a budget of $10\times 3000D=30000D$ target density evaluations per random seed is allocated for stochastic optimization. In the case of a noisy target, we further vary the observation noise level to study the noise sensitivity of the methods. For the Laplace approximation, we first find the MAP point by transforming the parameter space to unbounded if needed, via a nonlinear mapping. [^9] For the Laplace approximation to be valid, it is necessary to find the MAP estimate in the unbounded space subsequently used to compute the multivariate normal approximation; this is particularly important if the mode is close to the bounds. We subsequently compute the Hessian matrix via adaptive numerical differentiation (Brodtkorb and D’Errico, 2022). Finally, we compute the performance metrics as detailed below.
@@
    $\begin{split}\text{GsKL}(p,q)=&\frac{D_{\text{KL}}\left(\mathcal{N}[p]||\mathcal{N}[q]\right)}{2D}\\ &+\frac{D_{\text{KL}}(\mathcal{N}[q]||\mathcal{N}[p])}{2D},\end{split}$ (37)
@@ 4.3 Synthetic problems
$p(\mathbf{x})=$ $\log(\exp(\kappa x_{1}/r)/3+2\exp(-\kappa x_{1}/r)/3)$
$-\frac{1}{2}\left(\frac{r-1/\sqrt{2}}{0.1}\right)^{2},$ (38)
@@
(d) bbvi
Figure 4: Two Moons. The black density contours and points denote ground truth samples. The orange density contours and points represent the posterior samples from (a) vsbq; (b) nnr; (c) laplace; (d) bbvi with MoG ($K=50$). vsbq, nnr and bbvi with MoG ($K=50$) perfectly recover the bimodal density.
As shown in Figure 4 and Table 1, vsbq, nnr and bbvi with MoG($K=50$) reconstruct the bimodal target almost perfectly. By contrast, the Laplace approximation can only cope with unimodal posteriors and thus unsurprisingly fails in this case, despite its otherwise relative simplicity. bbvi with a diagonal Gaussian, a full-rank Gaussian, and a MoG($K=5$) also reveal inferior performance compared to the other methods.
@@
[…]
MoG ($K=5$) 0.44 $\scriptstyle{[0.28,0.79]}$ 0.21 $\scriptstyle{[0.13,0.28]}$ 5.7 $\scriptstyle{[0.16,8.7]}$
MoG ($K=50$) 0.0061 $\scriptstyle{[0.0021,0.015]}$ 0.022 $\scriptstyle{[0.018,0.031]}$ 0.00033 $\scriptstyle{[0.00012,0.00088]}$
laplace 0.43 0.19 8.1
@@
$\begin{split}p(\mathbf{x})\propto&~{}e^{\mathcal{R}(x_{1},x_{2})}e^{\mathcal{R}(x_{3},x_{4})}\mathcal{N}([x_{5},x_{6}];\mathbf{0},\mathbb{I})\\ &\cdot\mathcal{N}(\mathbf{x};\mathbf{0},3^{2}\mathbb{I}),\end{split}$ (39)
@@
As shown in Table 2, both vsbq and nnr approximate this complex posterior well, with vsbq performing slightly better than nnr in terms of metrics. laplace does not give a satisfactory approximation either, due to the heavily non-Gaussian nature of the underlying posterior. Among the bbvi methods, bbvi with MoG ($K=50$) achieves the best results. However, it still underperforms compared to vsbq, even when allocated a tenfold higher density evaluation budget. The poor metrics further indicate challenges in fitting a full-rank Gaussian using bbvi. A visualization of the approximate posteriors and the ground-truth posterior is provided in Supplementary Material C.6.
@@
Gaussian (diagonal) 1.2 $\scriptstyle{[1.1,1.3]}$ 0.23 $\scriptstyle{[0.23,0.23]}$ 0.55 $\scriptstyle{[0.54,0.57]}$
Gaussian (full-rank) 1.1e+03 $\scriptstyle{[3.8e+02,2.2e+03]}$ 0.66 $\scriptstyle{[0.64,0.75]}$ 3.2e+05 $\scriptstyle{[1.6e+04,1.4e+06]}$
MoG ($K=5$) 0.91 $\scriptstyle{[0.83,0.98]}$ 0.16 $\scriptstyle{[0.16,0.17]}$ 0.32 $\scriptstyle{[0.29,0.35]}$
MoG ($K=50$) 0.30 $\scriptstyle{[0.21,0.36]}$ 0.058 $\scriptstyle{[0.057,0.060]}$ 0.049 $\scriptstyle{[0.046,0.051]}$
laplace 1.3 0.24 0.91
@@ 4.4 Real-world models
We consider a popular Bayesian observer model of time perception from cognitive neuroscience (Jazayeri and Shadlen, 2010; Acerbi et al., 2012; Acerbi, 2020). The key premise of Bayesian observer modeling in perception is that the participant of a psychophysical experiment – the participant being the system being modeled – is herself performing Bayesian inference over the sensory stimuli, and employs Bayesian decision theory to report their perception (Pouget et al., 2013; Ma et al., 2023).
@@
In this specific sensorimotor timing experiment, in each trial human participants had to reproduce the time interval $\tau$ between a mouse click and a screen flash, with $\tau\sim\text{Uniform}$[0.6, 0.975] s (Acerbi et al., 2012). We assume participants had only access to a noisy sensory measurement $t_{\text{s}}\sim\mathcal{N}\left(\tau,w_{\text{s}}^{2}\tau^{2}\right)$, and their reproduced time $t_{\text{m}}$ was affected by motor noise, $t_{\text{m}}\sim\mathcal{N}\left(\tau_{\star},w_{m}^{2}\tau_{\star}^{2}\right)$, where $w_{\text{s}}$ and $w_{\text{m}}$ are Weber’s fractions, a psychophysical measure of perceptual and motor variability. We assume participants estimated $\tau_{\star}$ by combining their sensory likelihood with an approximate Gaussian prior over time intervals, $\mathcal{N}\left({\tau};{\mu_{\text{p}}},{\sigma^{2}_{\text{p}}}\right)$, and took the mean of the resulting Bayesian posterior. For each trial we also consider a probability $\lambda$ of a ‘lapse’ (e.g., a misclick) producing a response $t_{\text{m}}\sim\text{Uniform}$[0, 2] s. Model parameters are $\bm{\theta}=(w_{\text{s}},w_{\text{m}},\mu_{\text{p}},\sigma_{\text{p}},\lambda)$, so $D=5$. We infer the posterior of a representative participant using published data from Acerbi et al. (2012).
@@
As shown in Figure 5, vsbq consistently gives a good posterior approximation across different levels of log-likelihood evaluation noise. In this case, even with a large noise $\sigma_{\text{obs}}=7$, vsbq surpasses the performance of laplace. Note that laplace is computed based on noiseless target evaluations since it does not support noisy evaluations. In this problem, nnr is considerably less robust, with several failed runs, and performs generally worse compared to vsbq. Since the target posterior is close to a Gaussian, the bbvi methods exhibit similar performance to each other (as well as across noise levels), all fairly unsatisfactorily hovering above the desired metric thresholds. The insensitivity of bbvi performance to log-likelihood noise suggests that the variance of the score function estimator is the dominating factor behind the suboptimal results. bbvi with a Gaussian with full-rank covariance is excluded from the figure, as it yields poor results due to fitting challenges.
Plots of approximate posteriors and the ground-truth posteriors are visualized in Supplementary Material C.6.
@@
Figure 5: Bayesian timing model. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). vsbq performs well across all noise levels and can outperform (noiseless) laplace even under high log-likelihood noise, whereas nnr demonstrates less robustness with several failed runs. bbvi methods exhibit similar performance above the metric thresholds.
@@
Perceptual causal inference – inferring whether two sensory cues have the same common source – comprises a variety of models and tasks of major interest in computational and cognitive neuroscience (Körding et al., 2007; Cao et al., 2019; Shams and Beierholm, 2022). Here we consider a visuo-vestibular causal inference experiment representative of this class of models (Acerbi et al., 2018; Acerbi, 2020). In this experiment, participants were seated in a moving chair and asked to determine whether the direction of their movement ($s_{\text{vest}}$) corresponded to the direction of a looming visual field ($s_{\text{vis}}$) on a trial by trial basis. It is assumed that the participants can only access noisy sensory measurements, denoted as $z_{\text{vest}}\sim\mathcal{N}\left(s_{\text{vest}},\sigma^{2}_{\text{vest}}\right)$ for vestibular information and $z_{\text{vis}}\sim\mathcal{N}\left(s_{\text{vis}},\sigma^{2}_{\text{vis}}(c)\right)$ for visual information. Here, $\sigma_{\text{vest}}$ represents the vestibular noise, while $\sigma_{\text{vis}}(c)$ represents the visual noise, with $c$ being one of three distinct levels of visual coherence ($c_{\text{low}},c_{\text{med}},c_{\text{high}}$) used in the experiment. To model the participants’ responses, we use a heuristic ‘Fixed’ rule, which determines the source to be the same if the absolute difference between the visual and vestibular measurements is less than a threshold $\kappa$, i.e., $|z_{\text{vis}}-z_{\text{vest}}|<\kappa$. Additionally, the model incorporates a probability $\lambda$ of the participant providing a random response (Acerbi et al., 2018). The model parameters are $\theta=(\sigma_{\text{vis}}(c_{\text{low}}),\sigma_{\text{vis}}(c_{\text{med}}),\sigma_{\text{vis}}(c_{\text{high}}),\sigma_{\text{vest}},\lambda,\kappa)$, with a total of $D=6$ parameters. Here we fit data from participant S0 of Acerbi et al. (2018).
@@
Figure 6: Multisensory causal inference model. Median $\Delta$LML loss (left), MMTV (middle), and GsKL (right) as a function of the log-likelihood noise $\sigma_{\text{obs}}$ for the Bayesian Timing model. Shaded areas are 95% CI of the median and grey dash-dotted horizontal lines are the rule-of-thumb thresholds for good performance ($\Delta$LML=1, MMTV =0.2, GsKL =$1/8$). vsbq performs well across low to moderate noise levels and outperforms (noiseless) laplace even at high noise levels. nnr also demonstrates strong performance across all noise levels. bbvi methods perform similarly and are above the thresholds, except for the MoG($K=50$) at low noise levels.
@@
The performance metrics of all tested methods are plotted in Figure 6, as a function of log-likelihood observation noise. vsbq consistently outputs a good posterior approximation and only exceeds the desirable metrics thresholds at large observation noise. laplace works reasonably for the noiseless case, still worse than vsbq, and slightly above the thresholds for MMTV and GsKL. In this problem, nnr performs well, slightly outperforming vsbq in some metrics in a higher observation noise regime. bbvi with MoG ($K=50$) performs well for low noise cases and becomes similar to bbvi with a diagonal Gaussian or MoG ($K=5$) for high noise levels. bbvi with a full-rank Gaussian is excluded from the figure, as it yields poor results due to optimization challenges also in this case. We visualize the approximate and ground-truth posteriors in Supplementary Material C.6.